# Adaptive Width Neural Networks

### v4

@@ 3.1 Imposing a Soft Ordering on Neurons’ Importance
Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\boldsymbol{\lambda}\text{=}\boldsymbol{\nu},\boldsymbol{\theta}\text{=}\boldsymbol{\rho},x_{i})$ of Equation 4 as a modified MLP. Compared to a standard MLP, we use the variational parameters $\boldsymbol{\nu}$ that affect the truncation width at each hidden layer, whereas $\boldsymbol{\rho}$ are the weights. We choose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations. In other words, we impose a soft ordering of importance among neurons.
@@
It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$. This could lead to a degenerate situation where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and potentially wasting neurons. There are two strategies to address this undesirable effect. The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$. The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 or tanh activation [51]. We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
3.2 Rescaled Weight Initialization for Deep AWNN
Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWNN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme [25]. This affects convergence since gradients get close to zero and it becomes slow to train deep AWNN MLPs. We therefore derive a rescaled Kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@ 3.2 Rescaled Weight Initialization for Deep AWNN
Consider an MLP with activations as in Equation 10 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{j*}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
@@
Algorithm 1 AWNN Training Procedure
1:  Input: Dataset $\mathcal{D}$, initialized AWNN model $\mathcal{M}$ (Section 3.2)
2:  Output: Trained AWNN Model $\mathcal{M}$
@@
7:       $\text{loss}\leftarrow\text{ELBO}(\mathcal{M},\text{batch},\hat{y})$ // Eq. 9
@@
9:  function $\mathcal{M}$):
[…]
@@
12:       Use $D_{\ell}$ to update $\boldsymbol{\rho}_{\ell},\boldsymbol{\rho}_{\ell+1}$ // add/remove neurons
@@ 3.3 Future Directions and Limitations
MLPs are ubiquitous in modern deep architectures. They are used at the very end of CNNs, in each Transformer layer, and they process messages coming from neighbors in DGNs. Our experiments focus on MLPs to showcase AWNN’s broad applicability, but there are many other scenarios where one can apply AWNN’s principles. For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training.
From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle [56], which seeks maximally representative (i.e., performance) and compact representations (e.g., width). Finally, the analysis of different functions $f$ to soft-order neurons, for instance a power-law distribution or a sigmoid-like function, may have a different impact on convergence and performance and could be empirically investigated.
AWNN needs to change the network at training time with an associated overhead. Our current implementation has only a 3-4$\times$ overhead on the smallest datasets, the worst case scenario for our approach, where the inference cost is negligible compared to the network update. For more realistic tasks, we believe our method is advantageous compared to the cost of a grid-search or alternating training/updating approaches. In addition, further optimizations are possible: i) by updating the width every $M$ mini-batches rather than at every step; ii) by allocating a larger network and dynamically grow it only if needed; iii) by creating ad-hoc support of AWNN on deep learning libraries.
@@ 4 Experiments and Setup
The purpose of the empirical analysis is not to claim AWNN is generally better than the fixed-width baseline. Rather, we demonstrate how AWNN overcomes the problem of fixing the number of neurons by learning it end-to-end, thus reducing the amount of hyper-parameter configurations to test. As such, due to the nature of this work and similarly to Mitchell et al. [40], we use the remaining space to thoroughly study the behavior of AWNN, so that it becomes clear how to use it in practice. We first quantitatively verify that AWNN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWNN. Second, we check that AWNN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence. Third, we verify that convergence speed is not significantly altered by AWNN, so that the main limitation lies in the extra overhead for adapting the network at each training step. As a sanity check, we study conditions under which AWNN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much. Finally, we analyze other practical advantages of training a neural network under the AWNN framework: the ability to compress information during training or post training, and the resulting trade-offs. Further analyses are in the Appendix.
@@
We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWNN version, where we replace any fixed MLP with an adaptive one. First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard. A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets. Similarly, we consider a ResNet-20 [26] trained on 3 image classification tasks, namely MNIST [32], CIFAR10, and CIFAR100 [31], where data splits and preprocessing are taken from the original paper and AWNN is applied to the downstream classifier. In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset [67] with the same data split as MNIST. In the graph domain, we train a Graph Isomorphism Network [65] on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of Errica et al. [16]. Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWNN versions. On all these tasks, the metric of interest is the accuracy. Finally, for the textual domain we train a Transformer architecture [58] on the Multi30k English-German translation task [13], using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words. On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances after an early stopping strategy on the validation set. Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWNN versions in Appendix F and G, respectively [^3] Code to reproduce results is in the supplementary material.. We ran the experiments on a server with 64 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 48GB of memory.
@@ 5 Results
Table 1: Performances and total width of MLP layers for the fixed and AWNN versions of the various models used. The exact width chosen by model selection on the graph datasets is unknown since we report published results. “Linear" means the chosen downstream classifier is a linear model.
Fixed AWNN Width (Fixed) Width (AWNN)
@@
We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs. In terms of performance, we observe that AWNN is more stable or accurate than a fixed MLP on DoubleMoon, Spiral, SpiralHard, and PMNIST; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements. On the image datasets, performances of AWNN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge. There, AWNN learns a smaller total width compared to grid search.
Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWNN both in terms of average performance and stability of results; second, and akin to PMNIST, the total learned width is significantly higher than those tried in Xu et al. [65], Errica et al. [16], meaning a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN). This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task. Appendix H shows what happens when we retrain Fixed baselines using the total width as the width of each layer.
Finally, the results on the Multi30k show that the AWNN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss. This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models [9] is so high that cannot be afforded by most institutions, and it demands future investigations.
@@
Intuitively, one would expect that AWNN learned larger widths for more difficult tasks. This is indeed what happens on the tabular datasets (and image datasets, see Appendix I) where some tasks are clearly harder than others. Figure 2 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty. It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al. [3], Mhaskar et al. [37]. It also appears that convergence is not affected by the introduction of AWNN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
Figure 2: (Left) The learned width adapts to the increasing difficulty of the task, from the DoubleMoon to SpiralHard. (Right) AWNN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
@@
To support our argument that AWNN can reduce the time spent performing hyper-parameter selection, we check whether AWNN learns a consistent amount of neurons across different training runs and hyper-parameter choices. Figure 3 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection. Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width. Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix J shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1. Therefore, whenever possible, we recommend using bounded activations.
@@
So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width. We demonstrate the effect of an informative prior by performing a width-annealing experiment on the SpiralHard dataset. We set an uninformative $p(\boldsymbol{\theta})$ and ReLU6 nonlinearity. At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$. Figure 4 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance degradation. We hypothesize that the least important neurons carry negligible information, therefore they can be safely removed without drastic changes in the output of the model. This technique might be particularly useful to compress large models with billions of parameters.
Figure 4: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}$. The total width is reduced by more than 50% (left) while preserving accuracy (right). The inset plot refers to the loss term that AWNN tries to maximize.

### v5

@@ 3.1 Imposing a Soft Ordering on Neurons’ Importance
Now that the learning objective has been formalized, the missing ingredient is the definition of the neural network $p(y_{i}|\bm{\lambda}\text{=}\bm{\nu},\bm{\theta}\text{=}\bm{\rho},x_{i})$ of Equation 4 as a modified MLP. Compared to a standard MLP, we use the variational parameters $\bm{\nu}$ that affect the truncation width at each hidden layer, whereas $\bm{\rho}$ are the weights. We choose a monotonically decreasing function $f_{\ell}$, thus when a new neuron is added its relative importance is low and will not drastically impact the network output and hidden activations. In other words, we impose a soft ordering of importance among neurons.
@@
It is easy to see that, in theory, the optimization algorithm could rescale the weights of the next layer by a factor $1/f_{\ell}(j;\nu_{\ell})$ to compensate for the term $f_{\ell}(j;\nu_{\ell})$. This could lead to a degenerate situation where the activations of the first neurons are small relative to the others, thus breaking the soft-ordering and potentially wasting neurons. There are two strategies to address this undesirable effect. The first is to regularize the magnitude of the weights thanks to the prior $p(\theta_{\ell+1n})$, so that it may be difficult to compensate for the least important neurons that have a high $1/f_{\ell}(j)$. The second and less obvious strategy is to prevent the units’ activations of the current layer to compensate for high values by bounding their range, e.g., using a ReLU6 or tanh activation (Sandler et al., 2018). We apply both strategies to our experiments, although we noted that they do not seem strictly necessary in practice.
3.2 Rescaled Weight Initialization for Deep AWN
Rescaling the activations of hidden units using Equation 10 causes activations of deeper layers to quickly decay to zero for an AWN MLP with ReLU nonlinearity initialized using the well known Kaiming scheme (He et al., 2015). This affects convergence since gradients get close to zero and it becomes slow to train deep AWN MLPs. We therefore derive a rescaled Kaiming weight that guarantees that the variance of activation across layers is constant at initialization.
@@ 3.2 Rescaled Weight Initialization for Deep AWN
Consider an MLP with activations as in Equation 10 and ReLU nonlinearity. At initialization, given $\alpha_{j}^{\ell}=\sigma\left(\sum_{k=1}^{D_{\ell-1}}w^{\ell}_{jk}h^{\ell-1}_{k}\right),\mathrm{Var}[w_{jk}^{\ell}]=\frac{2}{\sum_{j=1}^{D_{\ell-1}}f^{2}_{\ell}(j)}\Rightarrow\mathrm{Var}[\alpha_{j}^{\ell}]\approx\mathrm{Var}[\alpha_{j}^{\ell-1}]$.
@@
Algorithm 1 AWN Training Procedure
1:Input: Dataset $\mathcal{D}$, initialized AWN model $\mathcal{M}$ (Section 3.2)
2:Output: Trained AWN Model $\mathcal{M}$
@@
[…]
@@
9:function update_width($\mathcal{M}$):
10:  for layer $\ell$ in $\mathcal{M}\text{.hidden\_layers}$ do
@@
12:   Use $D_{\ell}$ to update $\bm{\rho}_{\ell},\bm{\rho}_{\ell+1}$ // add/remove neurons
@@ 3.3 Future Directions and Limitations
MLPs are ubiquitous in modern deep architectures. They are used at the very end of CNNs, in each Transformer layer, and they process messages coming from neighbors in DGNs. Our experiments focus on MLPs to showcase AWN’s broad applicability, but there are many other scenarios where one can apply AWN’s principles. For instance, one could impose a soft ordering of importance on CNNs’ filters at each layer, therefore learning the number of filters during training. However, doing so would require a distinct formalization, which is why this lies beyond the scope of the present work.
From a more theoretical perspective, we believe one could draw connections between our technique and the Information Bottleneck principle (Tishby et al., 2000), which seeks maximally representative (i.e., performance) and compact representations (e.g., width). In addition, we could try to revisit the theoretical results of Caron et al. (2025) in the NTK regime to better understand convergence properties of adaptive-width neural networks. since both are based on asymmetric rescaling of activations.
@@ 4 Experiments and Setup
The purpose of the empirical analysis is not to claim AWN is generally better than the fixed-width baseline. Rather, we demonstrate how AWN overcomes the problem of fixing the number of neurons by learning it end-to-end, thus reducing the amount of hyper-parameter configurations to test. As such, due to the nature of this work and similarly to Mitchell et al. (2023), we use the remaining space to thoroughly study the behavior of AWN, so that it becomes clear how to use it in practice. We first quantitatively verify that AWN does not harm the performance compared to baseline models and compare the chosen width by means of grid-search model selection with the learned width of AWN. Second, we check that AWN chooses a larger width for harder tasks, which can be seen as increasing the hypotheses space until the neural network finds a good path to convergence. Third, we verify that convergence speed is not significantly altered by AWN, so that the main limitation lies in the extra overhead for adapting the network at each training step. As a sanity check, we study conditions under which AWN’s learned width does not seem to depend on starting hyper-parameters, so that their choice does not matter much. In addition, we analyze other practical advantages of training a neural network under the AWN framework: the ability to compress information during training or post training, and the resulting trade-offs. Finally, we analyze the impact of different families of functions $f_{\ell}(j;\nu_{\ell})$ compared to the exponential mainly used in this work. Further analyses are in the Appendix.
@@
We compare baselines that undergo proper hyper-parameter tuning (called “Fixed”) against its AWN version, where we replace any fixed MLP with an adaptive one. First, we train an MLP on 3 synthetic tabular tasks of increasing binary classification difficulty, namely a double moon, a spiral, and a double spiral that we call SpiralHard. A stratified hold-out split of 70% training/10% validation/20% test for risk assessment is chosen at random for these datasets. Then, we consider 3 larger real-world tabular tasks with higher feature dimensionality, namely pol, MiniBooNE, and credit card clients (Beyazit et al., 2023). Similarly, we consider a ResNet-20 (He et al., 2016) trained on 3 image classification tasks, namely MNIST (LeCun, 1998), CIFAR10, and CIFAR100 (Krizhevsky, 2009), where data splits and preprocessing are taken from the original paper and AWN is applied to the downstream classifier. In the sequential domain, we implemented a basic adaptive Recurrent Neural Network (RNN) and evaluated on the PMNIST dataset (Zenke et al., 2017) with the same data split as MNIST. In the graph domain, we train a Graph Isomorphism Network (Xu et al., 2019) on the NCI1 and REDDIT-B classification tasks, where topological information matters, using the same split and evaluation setup of Errica et al. (2020). Here, the first 1 hidden layer MLP as well as the one used in each graph convolutional layer are replaced by adaptive AWN versions. On all these tasks, the metric of interest is the accuracy. Finally, for the textual domain we train a Transformer architecture (Vaswani et al., 2017) on the Multi30k English-German translation task (Elliott et al., 2016), using a pretrained GPT-2 Tokenizer, and we evaluate the cross-entropy loss over the translated words. On tabular, image, and text-based tasks, an internal validation set (10%) for model selection is extracted from the union of outer training and validation sets, and the best configuration chosen according to the internal validation set is retrained 10 times on the outer train/validation/test splits, averaging test performances after an early stopping strategy on the validation set. Due to space reasons, we report datasets statistics and the hyper-parameter tried for the fixed and AWN versions in Appendix F and G, respectively [^2] Code to reproduce results is available at https://github.com/nec-research/Adaptive-Width-Neural-Networks.. We ran the experiments on a server with 64 cores, 1.5TB of RAM, and 4 NVIDIA A40 with 48GB of memory.
@@ 5 Results
Table 1: Performances and total width of MLP layers for the fixed and AWN versions of the various models used. The exact width chosen by model selection on the graph datasets is unknown since we report published results. “Linear" means the chosen downstream classifier is a linear model.
Fixed AWN Width (Fixed) Width (AWN)
@@
We begin by discussing the quantitative results of our experiments: Table 1 reports means and standard deviations across the 10 final training runs. In terms of performance, we observe that AWN is more stable or accurate than a fixed MLP on tabular and sequential datasets; all other things being equal, it seems that using more neurons and their soft ordering are the main contributing factors to these improvements. On the image datasets, performances of AWN are comparable to those of the fixed baseline but for CIFAR100, due to an unlucky run that did not converge. There, AWN learns a smaller total width compared to grid search.
Results on graph datasets are interesting in two respects: First, the performance on REDDIT-B is significantly improved by AWN both in terms of average performance and stability of results; second, and akin to PMNIST, the total learned width is significantly higher than those tried in Xu et al. (2019); Errica et al. (2020), meaning a biased choice of possible widths had a profound influence on risk estimation for DGN models (i.e., GIN). This result makes it evident that it is important to let the network decide how many neurons are necessary to solve the task. Appendix H shows what happens when we retrain Fixed baselines using the total width as the width of each layer.
Finally, the results on the Multi30k show that the AWN Transformer learns to use 200x parameters less than the fixed Transformer for the feed-forward networks, achieving a statistically comparable test loss. This result is appealing when read through the lenses of modern deep learning, as the power required by some neural networks such as Large Language Models (Brown et al., 2020) is so high that cannot be afforded by most institutions, and it demands future investigations.
@@
Intuitively, one would expect that AWN learned larger widths for more difficult tasks. This is indeed what happens on the tabular datasets (and image datasets, see Appendix I) where some tasks are clearly harder than others. Figure 2 (left) shows that, given the same starting width per layer, the learned number of neurons grows according to the task’s difficulty. It is also interesting to observe that the total width for a multi-layer MLP on SpiralHard is significantly lower than that achieved by a single-layer MLP, which is consistent with the circuit complexity theory arguments put forward in Bengio et al. (2006); Mhaskar et al. (2017). It also appears that convergence is not affected by the introduction of AWN, as investigated in Figure 2 (right), which was not obvious considering the parametrization constraints encouraged by the rescaling of neurons’ activations.
Figure 2: (Left) The learned width adapts to the increasing difficulty of the task, from the DoubleMoon to SpiralHard. (Right) AWN reaches perfect test accuracy with a comparable amount of epochs on DoubleMoon and Spiral, while it converges faster on SpiralHard.
@@
To support our argument that AWN can reduce the time spent performing hyper-parameter selection, we check whether AWN learns a consistent amount of neurons across different training runs and hyper-parameter choices. Figure 3 reports the impact of the batch size and starting width averaged across the different configurations tried during model selection. Smaller batch sizes cause more instability, but in the long run we observe convergence to a similar width. Convergence with respect to different rates holds, instead, for the bounded ReLU6 activation; Appendix K shows that unbounded activations may cause the network to converge more slowly to the same width, which is in accord with the considerations about counterbalancing the rescaling effect of Section 3.1. Therefore, whenever possible, we recommend using bounded activations.
@@
So far, we have used an uninformative prior $p(\lambda)$ over the neural networks’ width. We demonstrate the effect of an informative prior by performing a width-annealing experiment on the SpiralHard dataset. We set an uninformative $p(\bm{\theta})$ and ReLU6 nonlinearity. At epoch 1000, we introduce $p(\lambda_{\ell})=\mathcal{N}(\lambda_{\ell};0.05,1)$, and gradually anneal the standard deviation up to $0.1$ at epoch $2500$. Figure 4 shows that the width of the network reduces from approximately 800 neurons to 300 without any test performance degradation. We hypothesize that the least important neurons carry negligible information, therefore they can be safely removed without drastic changes in the output of the model. This technique might be particularly useful to compress large models with billions of parameters.
Figure 4: It is possible to regularize the width at training time by increasing the magnitude of the loss term $\log\frac{p(\bm{\nu})}{q(\bm{\nu})}$. The total width is reduced by more than 50% (left) while preserving accuracy (right). The inset plot refers to the loss term that AWN tries to maximize.
@@
Impact of Different Functions $f_{\ell}(j;\nu_{\ell})$
We conclude our main analyses with an ablation study on the use of different decreasing functions $f_{\ell}(j;\nu_{\ell})$, considering the SpiralHard dataset, to understand how different functions have different properties. We tested: i) a Power Law distribution $f_{\ell}(j;\nu_{\ell},\alpha)=Z(j+j_{sat})^{-\nu_{\ell}}$, with $Z$ being a normalizing term, initial $\nu_{\ell}\in[2.,3.]$ and low degree saturation $j_{sat}$ chosen between $[0.,2.]$; ii) a sigmoid-like function $f_{\ell}(j;\nu_{\ell},b)=1-\sigma(j-\nu_{\ell})$, where $\sigma$ is the sigmoidal activation, $\nu_{\ell}\in[128,256]$ is the transition point.
The results are reported in Table 2. When averaged over all possible hyper-parameter configurations, there is no significant deviation in performance between distribution, and there is always at least one configuration attaining the best performance. However, the Power Law distribution introduces more neurons than the exponential due to its long-tail. Instead, the Sigmoidal function tends to allocates less neurons, but the drawback is the loss of the ability to truncate the network after training, since the importance of neurons is very close to 1 before the transition point and the transition itself, for the chosen fixed parameters, is relatively sharp. The choice of the function family $f_{\ell}(j;\nu_{\ell})$ remains an interesting direction of further exploration, given that each family has different properties that might be more or less suitable depending on the use cases.
Table 2: Analysis of the impact of different families of importance functions on SpiralHard validation performance and learned width, averaged across different hyper-parameter configurations.
$f_{\ell}(j;\nu_{\ell})$ Mean Accuracy Max Accuracy Mean Total Width
Exponential 80.27 (19.9) 100.00 954.4 (1083.9)
Power Law 81.82 (16.6) 100.00 2952.4 (3371.6)
Sigmoidal 76.85 (18.3) 100.00 426.8 (268.4)